OpenStudy (anonymous):

Solving equation help please? Solve the equation in the real number system: x^5 - 4x^4 + 4x^3 + 2x^2 - 5x + 2 = 0 (A) x = -5, -4, 1, 2 (B) x = -1, 1, 2 (C) x = -5, -4, 1, 2, 4 (D) x = -2, -1, 1

OpenStudy (danjs):

(x^3 - 3x +3x - 1)*(x^2 - x - 2) = 0

OpenStudy (danjs):

(x-1)^3 * (x + 1)*(x - 2) = 0

OpenStudy (danjs):

(x-1)=0 or (x+1)=0 or (x-2) = 0

OpenStudy (danjs):

x=1 or x=-1 or x=2

OpenStudy (anonymous):

thank you

OpenStudy (danjs):

yw